# Regist3R: Incremental Registration with Stereo Foundation Model

### v1

@@ Abstract
1 Introduction
@@
As the cornerstone of multi-view 3D reconstruction, Structure from Motion (SfM) [1] reconstructs 3D structures and camera poses from unordered 2D images, but struggles with feature matching noise, scale drift, and scalability. Although global SfM [2, 3, 4] optimizes all camera parameters simultaneously, its dependency on reliable initial pairwise geometries limits its effectiveness in scenes with sparse features. Incremental SfM [5, 6, 7] mitigates these issues through sequential image registration and iterative refinement, proving to be more resilient in unstructured environments where global methods often fail. This adaptability explains its prevalent use in real-world applications with imperfect data conditions.
@@
Traditional SfM pipelines rely on sequential modules—feature extraction, matching, relative pose estimation, and triangulation [8, 9]—a fragmented paradigm susceptible to cumulative error propagation. Recent learning-based methods [10, 11, 12] attempt to replace these handcrafted components with end-to-end formulations. While VGGSfM [10] introduces differentiable submodules and Detector-free SfM [13] employs learned feature matching, both retain the classical pipeline structure. More radical approaches like FlowMap [14] and Ace-Zero [11] abandon modular designs entirely, directly optimizing geometry via per-scene gradient descent with geometry regressor networks. However, such methods remain constrained to scenarios with high image overlap and stable illumination.
@@
In this landscape, DUSt3R [12] emerges as a breakthrough by predicting globally consistent pointmaps in a unified coordinate system from two views, demonstrating exceptional geometric coherence without incremental error accumulation. Building on this success, recent extensions [15, 16, 17, 18, 19, 20, 21, 22] explore multi-view generalizations of DUSt3R’s framework, which demonstrate divergent optimization paradigms. Post-optimization approaches, exemplified by MASt3R-SfM [16] and Light3R-SfM [20], iteratively reconstruct scenes from pairwise matchings: MASt3R-SfM leverages pointmaps and matches from MASt3R for bundle adjustment, while Light3R-SfM accelerates convergence via direct Procrustes alignment [23]. However, these methods fail to establish fully differentiable reconstruction pipelines driven by Transformer inference. Alternatively, another research direction [17, 18] explores direct multi-view geometry prediction through scaled Transformer architectures. By integrating cross-attention mechanisms, these models enforce coordinate consistency across views during single-forward passes. Yet their scalability remains fundamentally constrained by the quadratic complexity growth of cross-attention layers with increasing view numbers. These limitations expose a critical gap in current learning-based SfM systems—achieving both optimization-free and scalable multi-view generalization.
@@
To address this challenge, we present Regist3R, a stereo foundational model for incremental registration. Regist3R is a transformer-based [24] network that regresses the pointmap of an unregistered view in the world coordinate system by leveraging a given reference view along with its world coordinate pointmap. An intuitive formulation and illustration compared to DUSt3R is shown as Eq. 1 and Fig. 1,
$I^{1}+I^{2}$ $\xrightarrow{DUSt3R}X^{1,1}+X^{2,1}$ (1)
@@
Drifting is one of the most critical problems for incremental reconstruction [25], and time-consuming bundle adjustment (BA) is adopted by most of traditional pipelines [5]. In the absence of BA for inference-based models, we enhance Regist3R’s noise resistance to mitigate drift. Specifically, in addition to feeding the model ground truth pointmap as input, we use a chain training strategy to estimate target structure based on inaccurate pointmap.
For efficient inference, we build a minimum spanning tree (MST) based on the view similarity, and perform pairwise inference only between the parents and children. This takes only $N-1$ times inference for a collection containing $N$ images. As the drifting happens when the tree goes deep, we adopt a tree compression trick to reduce the length of reconstruction chain.
@@
2 Related Works
2.1 Structure-from-Motion
@@
Structure-from-Motion (SfM) [1] is a pivotal technique in computer vision for reconstructing 3D scenes from 2D images. Traditional SfM pipelines predominantly employ two approaches: global and incremental methods.
@@
Global SfM [26, 2, 4, 27, 28] addresses the entire dataset simultaneously, estimating camera positions and 3D structures in a unified optimization process. While this approach offers efficiency, it often faces challenges related to scalability and precision. The computational demands can be substantial, especially with large datasets, and maintaining accuracy across extensive reconstructions remains a concern. Recent advancements, such as the GLOMAP [29] and XM [30], aim to enhance global SfM by balancing efficiency with improved accuracy, making it more viable for complex modeling tasks.
@@
Incremental SfM [31, 7, 5, 32], on the other hand, reconstructs scenes by progressively adding images, starting from an initial pair and iteratively integrating new ones. This method is renowned for its accuracy and robustness, particularly in well-textured environments. However, its time complexity is often considered $O(n^{4})$ concerning the number of images, leading to inefficiencies as dataset sizes grow. Hybrid SfM [33, 34, 35] approaches attempt to combine both for balance between efficiency and scalability.
@@
Traditional SfM methods, particularly those based on optimization, often encounter limitations in both efficiency and accuracy. The computational burden of global optimization can be prohibitive, and the incremental nature of certain methods may not scale well with increasing data volume. Although learning-based methods [9, 36, 37, 38, 39, 40, 41, 42, 43] replace some steps, the cumulative error of the pipeline still limits its scalability. In contrast, our model directly learns the incremental registration of SfM, eliminating the cumulative error caused by the pipeline. The inference-only mode eliminates the need for optimization and improves registration efficiency.
2.2 Reconstruction in Large Model Era
@@
Instead of replacing some steps of the SfM pipeline, recent works have proposed to learn the reconstruction task in an end-to-end manner [44, 10, 11]. VGGSfM [10] makes individual SfM components learnable forming a fully differentiable SfM framework. ACEZero[11] proposes to incrementally optimize scene coordinate regression and camera refinement networks to minimize reprojection errors. Similarly, FlowMap optimizes the per-scene depth estimation network with offline optical flow and point tracking supervision. Yet the performance is limited when the image set shares low visual overlapping.
@@
The recently proposed Dense and Unconstrained Stereo 3D Reconstruction model DUSt3R [12] brings a new paradigm to 3D reconstruction. It proposed a novel approach for two-view reconstruction via direct pointmap regression from a pair of RGB images, and the pointmaps of two input views share the same coordinate system. The improved version MASt3R [15] further extends DUSt3R with image matching. When handling multi-view reconstruction, the exhaustive pairwise inference constrains the scalability of DUSt3R. MASt3R-SfM [16] incorporates image retrieval and builds an optimization pipeline on MASt3R matches. However, the optimization process is time-consuming and the precision relies on matching accuracy. Multiple works extend DUSt3R to multi-view scenarios. Analogous to traditional approaches, we categorize them into global and incremental. Global approaches extend the model to accept multi-view images and use wide cross-attention for information sharing. Fast3R [17] uses random index embedding to improve model generalizability. MV-DUSt3R [18] uses cross-reference-view blocks to make it robust to reference view selection. VGGT [19] proposes to predict multiple non-orthogonal variables, like camera pose, depthmap, and pointmap, meanwhile using alternative attention for global alignment. Light3R-SfM [20] adopts a latent global alignment module to reduce the cost of global attention. However, the training cost is high and the scalability is limited by the model generalizability. Incremental approaches estimate the 3D structure of the current view based on previous reconstructed structures. Spann3R [22] maintains a memory bank for sequential reconstruction, but the scenario is limited to sequential images, and the generalizability of implicit memory bank representation of history frames is questionable. MUSt3R [21] proposes a similar implicit memory mechanism, therefore, its application scenarios are still limited to online reconstruction of sequential images.
@@
Figure 1: Comparison between Regist3R and DUSt3R. (Top) Two output pointmaps of DUSt3R share the same coordinate system, (Bottom) while the input and output pointmaps of Regist3R share the same coordinate system. The accompanying confidences are omitted.
3 Approach
The two-view foundation model DUSt3R [12] accepts two images in the same coordinate system and outputs their respective point maps. However, when extending to multiple views, it requires exhaustively traversing all pairs and performing post-optimization. Subsequent multi-view extensions [18, 17, 19] are still limited to a relatively small scale of views. Therefore, a crucial question is how to design a foundational model that achieves true scalability. In this section, we analyze and address this issue from the perspectives of problem formulation, model architecture, training procedure, and inference strategy.
3.1 Problem Formulation
@@ 3.2 Model Architecture
for $l=1,...,B$ for decoder with $B$ blocks and initialized with encoder tokens $G_{0}^{i}:=F^{i}$ and $G_{0}^{j}:=F^{j}$. The target decoder weights are initialized with DUSt3R’s first decoder and the reference decoder weights are optimized from scratch.
@@
$X^{j},C^{j}=\text{Head}(G^{j}_{0},...,G^{j}_{B}).$ (4)
As proved by DUSt3R, two-stream network architecture achieves impressive performance on relative pointmap prediction. Meanwhile, multi-view network architectures [17, 19] suffer from long sequence cross-attention and generalizability on more views than training set. Therefore, we keep the two-stream architecture for Regist3R, and extend it to multi-view scenarios by autoregression. In Sec. 3.4, we will demonstrate the advantages of Regist3R’s two-stream architecture in long-sequence inference.
3.3 Training Procedure
3.3.1 Pointmap Preperation
@@
$\mathcal{N}_{i}(X)=\frac{X-\mu(\bar{X}^{i})}{z(\bar{X}^{i})},\quad z(\bar{X}^{i})=\text{norm}\left(\bar{X}^{i}-\mu(\bar{X}^{i})\right).$ (5)
@@
$X^{j},C^{j}=f_{\theta}\left[I^{i}\|\mathcal{N}_{i}\left(R\bar{X}^{i}\right),I^{j}\right]$ (6)
@@
$\mathcal{L}_{\text{regr}}(i,j)$ $=\left\|\mathcal{N}_{i}\left(R\bar{X}^{j}\right)-X^{j}\right\|$ (7)
@@ 3.3.2 Autoregressive Training
During incremental registration, the precision of the reference pointmap is not guaranteed and the model should be robust to noise. To achieve this, we adopt a chain training strategy, where the model is trained to estimate the target structure based on the inaccurate pointmap and the associated confidence. Sepecifically, we sample a chain of images and ground-truth pointmaps $(I^{1},\bar{X}^{1}),...,(I^{N},\bar{X}^{N})$. For each step, the previously predicted pointmap and confidence are used as input. As the confidence value is activated by $exp$, we normalize it with log-sigmoid, and the confidence of ground truth is all one. Given the output pointmap of previous step $X^{n},C^{n}(n=1,...,N-1)$, and the image pair $I^{n},I^{n+1}$, the next step pointmap $X^{n+1},C^{n+1}$ is estimated by:
$X^{n+1},C^{n+1}=f_{\theta}\left[I^{n}\|\mathcal{N}_{n}\mathcal{N}_{n-1}^{-1}X^{n}\|\sigma\circ\log(C^{n}),I^{n+1}\right]$ (8)
@@
The patch embedding of reference encoder is extended to 7 channels (RGB-XYZ-C) to fit confidence input. Different from Spann3R [22] which transfers implicit features across steps and thus requires cross-step optimization, our Regist3R uses pointmap and confidence to transfer information, so each step is trained independently and there is no need to transfer gradients between steps. A training procedure with a chain length of 3 is shown in Fig. 2.
Figure 2: Autoregressive training strategy.
3.4 Inference Strategy

### v2

@@ Abstract.
Multi-view 3D Reconstruction, Structure-from-Motion, 3D Foundation Model.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746027.3755466conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Reconstructionccs: Computing methodologies Matching
Figure 1. Qualitative examples of Regist3R, covering indoor, outdoor, and aerial scenes, demonstrating its generalizability.
1. Introduction
@@
As the cornerstone of multi-view 3D reconstruction, Structure from Motion (SfM) (Agarwal et al., 2011) reconstructs 3D structures and camera poses from unordered 2D images, but struggles with feature matching noise, scale drift, and scalability. Although global SfM (Moulon et al., 2013; Cai et al., 2021; Wilson and Snavely, 2014) optimizes all camera parameters simultaneously, its dependency on reliable initial pairwise geometries limits its effectiveness in scenes with sparse features. Incremental SfM (Schonberger and Frahm, 2016; Snavely et al., 2006; Wu, 2013) mitigates these issues through sequential image registration and iterative refinement, proving to be more resilient in unstructured environments where global methods often fail. This adaptability explains its prevalent use in real-world applications with imperfect data conditions.
@@
Traditional SfM pipelines rely on sequential modules—feature extraction, matching, relative pose estimation, and triangulation (Lowe, 2004; Sarlin et al., 2020)—a fragmented paradigm susceptible to cumulative error propagation. Recent learning-based methods (Wang et al., 2024a; Brachmann et al., 2024; Wang et al., 2024b) attempt to replace these handcrafted components with end-to-end formulations. While VGGSfM (Wang et al., 2024a) introduces differentiable submodules and Detector-free SfM (He et al., 2024) employs learned feature matching, both retain the classical pipeline structure. More radical approaches like FlowMap (Smith et al., 2024) and Ace-Zero (Brachmann et al., 2024) abandon modular designs entirely, directly optimizing geometry via per-scene gradient descent with geometry regressor networks. However, such methods remain constrained to scenarios with high image overlap and stable illumination.
@@
In this landscape, DUSt3R (Wang et al., 2024b) emerges as a breakthrough by predicting globally consistent pointmaps in a unified coordinate system from two views, demonstrating exceptional geometric coherence without incremental error accumulation. Building on this success, recent extensions (Leroy et al., 2024; Duisterhof et al., 2024; Yang et al., 2025; Tang et al., 2024; Wang et al., 2025; Elflein et al., 2025; Cabon et al., 2025; Wang and Agapito, 2024) explore multi-view generalizations of DUSt3R’s framework, which demonstrate divergent optimization paradigms. Post-optimization approaches, exemplified by MASt3R-SfM (Duisterhof et al., 2024) and Light3R-SfM (Elflein et al., 2025), iteratively reconstruct scenes from pairwise matchings: MASt3R-SfM leverages pointmaps and matches from MASt3R for bundle adjustment, while Light3R-SfM accelerates convergence via direct Procrustes alignment (Umeyama, 1991). However, these methods fail to establish fully differentiable reconstruction pipelines driven by Transformer inference. Alternatively, another research direction (Yang et al., 2025; Tang et al., 2024) explores direct multi-view geometry prediction through scaled Transformer architectures. By integrating cross-attention mechanisms, these models enforce coordinate consistency across views during single-forward passes. Yet their scalability remains fundamentally constrained by the quadratic complexity growth of cross-attention layers with increasing view numbers. These limitations expose a critical gap in current learning-based SfM systems—achieving both optimization-free and scalable multi-view generalization.
@@
To address this challenge, we present Regist3R, a stereo foundational model for incremental registration. Regist3R is a transformer-based (Vaswani et al., 2017) network that regresses the pointmap of an unregistered view in the world coordinate system by leveraging a given reference view along with its world coordinate pointmap. An intuitive formulation and illustration compared to DUSt3R is shown as Eq. 1 and Fig. 2,
(1) $I^{1}+I^{2}$ $\xrightarrow{DUSt3R}X^{1,1}+X^{2,1}$
@@
Drifting is one of the most critical problems for incremental reconstruction (Holynski et al., 2020), and time-consuming bundle adjustment (BA) is adopted by most of traditional pipelines (Schonberger and Frahm, 2016). In the absence of BA for inference-based models, we enhance Regist3R’s noise resistance to mitigate drift. Specifically, in addition to feeding the model ground truth pointmap as input, we use a chain training strategy to estimate target structure based on inaccurate pointmap.
For efficient inference, we build a minimum spanning tree (MST) based on the view similarity, and perform pairwise inference only between the parents and children. This takes only N1N-1 times inference for a collection containing $N$ images. As the drifting happens when the tree goes deep, we adopt a tree compression trick to reduce the length of reconstruction chain.
@@
2. Related Works
2.1. Structure-from-Motion
@@
Structure-from-Motion (SfM) (Agarwal et al., 2011) is a pivotal technique in computer vision for reconstructing 3D scenes from 2D images. Traditional SfM pipelines predominantly employ two approaches: global and incremental methods.
@@
Global SfM (Jiang et al., 2013; Moulon et al., 2013; Wilson and Snavely, 2014; Ozyesil and Singer, 2015; Moulon et al., 2016) addresses the entire dataset simultaneously, estimating camera positions and 3D structures in a unified optimization process. While this approach offers efficiency, it often faces challenges related to scalability and precision. The computational demands can be substantial, especially with large datasets, and maintaining accuracy across extensive reconstructions remains a concern. Recent advancements, such as the GLOMAP (Pan et al., 2024) and XM (Han and Yang, 2025), aim to enhance global SfM by balancing efficiency with improved accuracy, making it more viable for complex modeling tasks.
@@
Incremental SfM (Snavely et al., 2008; Wu, 2013; Schonberger and Frahm, 2016; Mur-Artal and Tardós, 2017), on the other hand, reconstructs scenes by progressively adding images, starting from an initial pair and iteratively integrating new ones. This method is renowned for its accuracy and robustness, particularly in well-textured environments. However, its time complexity is often considered $O(n^{4})$ concerning the number of images, leading to inefficiencies as dataset sizes grow. Hybrid SfM (Cui et al., 2017; Zhu et al., 2017; Ye et al., 2024b, a) approaches attempt to combine both for balance between efficiency and scalability.
@@
Traditional SfM methods, particularly those based on optimization, often encounter limitations in both efficiency and accuracy. The computational burden of global optimization can be prohibitive, and the incremental nature of certain methods may not scale well with increasing data volume. Although learning-based methods (Sarlin et al., 2020; Lindenberger et al., 2021; Dusmanu et al., 2019; Gleize et al., 2023; Teed and Deng, 2021; Yi et al., 2016; Wang et al., 2023; Li et al., 2025; Zhang et al., 2024) replace some steps, the cumulative error of the pipeline still limits its scalability. In contrast, our model directly learns the incremental registration of SfM, eliminating the cumulative error caused by the pipeline. The inference-only mode eliminates the need for optimization and improves registration efficiency.
2.2. Reconstruction in Large Model Era
@@
Instead of replacing some steps of the SfM pipeline, recent works have proposed to learn the reconstruction task in an end-to-end manner (Zhao et al., 2022; Wang et al., 2024a; Brachmann et al., 2024). VGGSfM (Wang et al., 2024a) makes individual SfM components learnable forming a fully differentiable SfM framework. ACEZero(Brachmann et al., 2024) proposes to incrementally optimize scene coordinate regression and camera refinement networks to minimize reprojection errors. Similarly, FlowMap optimizes the per-scene depth estimation network with offline optical flow and point tracking supervision. Yet the performance is limited when the image set shares low visual overlapping.
@@
The recently proposed Dense and Unconstrained Stereo 3D Reconstruction model DUSt3R (Wang et al., 2024b) brings a new paradigm to 3D reconstruction. It proposed a novel approach for two-view reconstruction via direct pointmap regression from a pair of RGB images, and the pointmaps of two input views share the same coordinate system. The improved version MASt3R (Leroy et al., 2024) further extends DUSt3R with image matching. When handling multi-view reconstruction, the exhaustive pairwise inference constrains the scalability of DUSt3R. MASt3R-SfM (Duisterhof et al., 2024) incorporates image retrieval and builds an optimization pipeline on MASt3R matches. However, the optimization process is time-consuming and the precision relies on matching accuracy. Multiple works extend DUSt3R to multi-view scenarios. Analogous to traditional approaches, we categorize them into global and incremental. Global approaches extend the model to accept multi-view images and use wide cross-attention for information sharing. Fast3R (Yang et al., 2025) uses random index embedding to improve model generalizability. MV-DUSt3R (Tang et al., 2024) uses cross-reference-view blocks to make it robust to reference view selection. VGGT (Wang et al., 2025) proposes to predict multiple non-orthogonal variables, like camera pose, depthmap, and pointmap, meanwhile using alternative attention for global alignment. Light3R-SfM (Elflein et al., 2025) adopts a latent global alignment module to reduce the cost of global attention. However, the training cost is high and the scalability is limited by the model generalizability. Incremental approaches estimate the 3D structure of the current view based on previous reconstructed structures. Spann3R (Wang and Agapito, 2024) maintains a memory bank for sequential reconstruction, but the scenario is limited to sequential images, and the generalizability of implicit memory bank representation of history frames is questionable. MUSt3R (Cabon et al., 2025) proposes a similar implicit memory mechanism, therefore, its application scenarios are still limited to online reconstruction of sequential images.
@@
Figure 2. Comparison between Regist3R and DUSt3R. (Top) Two output pointmaps of DUSt3R share the same coordinate system, (Bottom) while the input and output pointmaps of Regist3R share the same coordinate system. The accompanying confidences are omitted.
3. Approach
The two-view foundation model DUSt3R (Wang et al., 2024b) accepts two images in the same coordinate system and outputs their respective point maps. However, when extending to multiple views, it requires exhaustively traversing all pairs and performing post-optimization. Subsequent multi-view extensions (Tang et al., 2024; Yang et al., 2025; Wang et al., 2025) are still limited to a relatively small scale of views. Therefore, a crucial question is how to design a foundational model that achieves true scalability. In this section, we analyze and address this issue from the perspectives of problem formulation, model architecture, training procedure, and inference strategy.
3.1. Problem Formulation
@@ 3.2. Model Architecture
for l=1,,Bl=1,...,B for decoder with $B$ blocks and initialized with encoder tokens $G_{0}^{i}:=F^{i}$ and $G_{0}^{j}:=F^{j}$. The target decoder weights are initialized with DUSt3R’s first decoder and the reference decoder weights are optimized from scratch.
@@
(4) $X^{j},C^{j}=\text{Head}(G^{j}_{0},...,G^{j}_{B}).$
As proved by DUSt3R, two-stream network architecture achieves impressive performance on relative pointmap prediction. Meanwhile, multi-view network architectures (Yang et al., 2025; Wang et al., 2025) suffer from long sequence cross-attention and generalizability on more views than training set. Therefore, we keep the two-stream architecture for Regist3R, and extend it to multi-view scenarios by autoregression. In Sec. 3.4, we will demonstrate the advantages of Regist3R’s two-stream architecture in long-sequence inference.
3.3. Training Procedure
3.3.1. Pointmap Preperation
@@
(5) $\mathcal{N}_{i}(X)=\frac{X-\mu(\bar{X}^{i})}{z(\bar{X}^{i})},\quad z(\bar{X}^{i})=\text{norm}\left(\bar{X}^{i}-\mu(\bar{X}^{i})\right).$
@@
(6) $X^{j},C^{j}=f_{\theta}\left[I^{i}\|\mathcal{N}_{i}\left(R\bar{X}^{i}\right),I^{j}\right]$
@@
(7) $\mathcal{L}_{\text{regr}}(i,j)$ $=\left\|\mathcal{N}_{i}\left(R\bar{X}^{j}\right)-X^{j}\right\|$
@@ 3.3.2. Autoregressive Training
During incremental registration, the precision of the reference pointmap is not guaranteed and the model should be robust to noise. To achieve this, we adopt a chain training strategy, where the model is trained to estimate the target structure based on the inaccurate pointmap and the associated confidence. Sepecifically, we sample a chain of images and ground-truth pointmaps $(I^{1},\bar{X}^{1}),...,(I^{N},\bar{X}^{N})$. For each step, the previously predicted pointmap and confidence are used as input. As the confidence value is activated by expexp, we normalize it with log-sigmoid, and the confidence of ground truth is all one. Given the output pointmap of previous step $X^{n},C^{n}(n=1,...,N-1)$, and the image pair $I^{n},I^{n+1}$, the next step pointmap $X^{n+1},C^{n+1}$ is estimated by:
(8) $X^{n+1},C^{n+1}=f_{\theta}\left[I^{n}\|\mathcal{N}_{n}\mathcal{N}_{n-1}^{-1}X^{n}\|\sigma\circ\log(C^{n}),I^{n+1}\right]$
@@
The patch embedding of reference encoder is extended to 7 channels (RGB-XYZ-C) to fit confidence input. Different from Spann3R (Wang and Agapito, 2024) which transfers implicit features across steps and thus requires cross-step optimization, our Regist3R uses pointmap and confidence to transfer information, so each step is trained independently and there is no need to transfer gradients between steps. A training procedure with a chain length of 3 is shown in Fig. 3.
Figure 3. Autoregressive training strategy.
3.4. Inference Strategy